What is the value of c, a value from the t distribution for a sample size of 14,such that the probability of being greater than c 1%
Answer #1

Solution,

Given that,

sample size = n = 14

degrees of freedom = n - 1 = 14 - 1 = 13

P( t > c ) = 0.01

= 1 - P( t < c ) = 0.01

= P( t < c ) = 1 - 0.01

= P( t < c ) = 0.99

= P( t < 2.650) = 0.99

c = 2.650